OpenStudy (anonymous):

Simplify seventeen s plus three t minus nine minus five t plus nine minus twelve s. fourteen s minus two t minus nine three s t five s minus two t fourteen s minus t

OpenStudy (texaschic101):

17s + 3t - 9 - 5t + 9 - 12s -- is this your problem ?

OpenStudy (anonymous):

yea thats it

OpenStudy (anonymous):

i think the aswer is 3st

OpenStudy (anonymous):

but i forgot to simplify so that cant be right

OpenStudy (texaschic101):

we will combine like terms... 17s - 12s = 5s 3t - 5t = -2t -9 + 9 = 0 put them together and we get : 5s - 2t

OpenStudy (anonymous):

okay wow thanks

OpenStudy (texaschic101):

anytime....happy to help :)

OpenStudy (anonymous):

that doe make alot of since now that u put it like that

OpenStudy (anonymous):

does i mena

OpenStudy (texaschic101):

whatever makes it easier to do is a good thing...lol